Quiz Three Name Directions: Put your answers on the actual quiz where yousee the words PLACEFINAL Staple yourwork to the rear of the stapled quiz and make sure it is organized for analysis of your computations. Lastly, type (or print clearly) and work in which your interpretation is required Dr. Baker believes that virtual reality is one of the biggest technological accomplishments in the 21 century. He creates a statistical program implemented through virtual reality and to testits effectiveness he gives his students a standardized calculus test in weekone of the semester then gives each student a pair of virtual reality googles with his eight week statistical program. He now gives his students another standardized calculus test in week tenof the semester. Did the virtual reality program prove successful at ? .05? 1) He PLACE FINAL ANSWER HERE Week 8 58 Week 1 42 845 50 23 61 50 45 85 90 26 20 35 50 42 48 58 60 PLACE FINAL ANSWERS HERE 77, o 8, at2 .05? 2) what is the confidence interval for ? -218,02-35, at2 : .05? 3) what is the confidence interval for 4) What is the confidence interval for ?-357, ?2-28, at 2-05?
Answer #1

Result:

1).

week1

week2

42

58

45

50

23

61

50

55

45

58

85

90

26

30

20

35

50

42

48

60

Null hypothesisH0·Ha = 0 Alternative hypothesis . Hi ·Ha < 0

Lower tail test

Sd

Paired t Test

Data

Hypothesized Mean Difference

0

Level of significance

0.05

Intermediate Calculations

Sample Size

10

DBar

-10.5000

Degrees of Freedom

9

SD

11.9745

Standard Error

3.7867

t Test Statistic

-2.7729

Lower-Tail Test

Lower Critical Value

-1.8331

p-Value

0.0108

Reject the null hypothesis

Rejection Region: Reject Ho if t < -1.8331

Calculated t = -2.7729 falls in the rejection region

The null hypothesis is rejected.

We conclude that virtual reality program proved successful.

2).

Z value at 0.05 level = 1.96

Lower limit = 77-1.96*8 =61.32

Upper limit = 77+1.96*8 = 92.68

3).

Lower limit = 218-1.96*sqrt(35) = 206.4045

Upper limit = 218+1.96*sqrt(35) = 229.5955

4).

Lower limit = 357-1.96*sqrt(28) = 346.6287

upper limit = 357+1.96*sqrt(28) = 367.3713